OpenStudy (anonymous):

1.Create an equation that results in at least one extraneous solution. Work through your equation, justify each step, and explain how the solution is extraneous. 2.Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x². Help quick I will award a medal

OpenStudy (anonymous):

@Mertsj help please need this to pass

OpenStudy (mertsj):

\[\sqrt{x-3}=-5\]

OpenStudy (anonymous):

is this to the first one

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

can you help me with number 2 please

OpenStudy (anonymous):

@Mertsj do you think you could help me with number 2 and explain please

OpenStudy (mertsj):

Simplify this: \[\sqrt[3]{x^6}\]

OpenStudy (anonymous):

I have no idea how to do this

OpenStudy (mertsj):

I don't know what to say. You need to know the laws of exponents to do these problems and you surely have been studying them for some time to get as far as radicals.